OpenStudy (anonymous):

What is the Domain and range of a circle with the equation of x^2 + y^2 = 4, and a radius of 2?

OpenStudy (zehanz):

You already know the radius of the circle (2). Do you also know the coordinates of the center?

OpenStudy (zehanz):

The equation of a circle with center (a, b) and radius r is: \((x-a)^2+(y-b)^2=r^2\). Because you have \(x^2+y^2=2^2\), you can see that the origin is the center. Now think: if the radius is 2, then what are the minimum and maximum values x AND y can be?